Circular No. 7148

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 F1 (CATALINA) On Apr. 17, T. B. Spahr, Lunar and Planetary Laboratory, reported

the automatic discovery of an object of unusual motion and stellar appearance in the course of the Catalina Sky Survey (0.41-m Schmidt + CCD):

1999 UT  α_{2000}  δ_{2000}  m_{2}
Mar. 23.31633  13^{h}22^{m}09.^{s}56  −8^{o}27′59′′.9  18.3

Spahr obtained follow-up data on Apr. 16 and 17. Computations by B. G. Marsden suggested that the object was a long-period comet in a highlyinclined orbit, yielding an identification in Mar. 13 Spacewatch data. Further observations (given on MPEC 1999-H09) produce the orbital elements below. CCD images (660 s total exposure) obtained with the Catalina 1.5m reflector by J. Bialozynski, D. Dietrich, C. Greenberg, E. Hooper, D. McBee, D. McCarthy, J. Pici, G. Rudnick, and C. Vedeler, and co-added by C. W. Hergenrother, show a faint coma of diameter 8′′–10′′.

T = 2002 Feb. 13.883 TT  ω = 254.^{o}994
Ω = 20.036 }2000.0
q = 5.80301 AU  i = 92.160
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Apr. 12 13^{h}12.^{m}07 − 8^{o}24.^{′}1 8.140 9.142 178.^{o}0 0.^{o}2 18.9
22 13 06.80 − 8 21.4 8.111 9.093 167.1 1.4 19.0
May 2 13 01.71 − 8 19.1 8.114 9.045 156.2 2.6 19.1
12 12 56.94 − 8 18.3 8.147 8.996 145.4 3.7 19.2
22 12 52.67 − 8 19.6 8.206 8.948 134.8 4.6 19.2

SUPERNOVA 1999bv IN MCG +10-25-14 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag ∼ 18.1) on an unfiltered CCD frame taken on Apr. 19.446 UT, utilizing the Tenagra I 0.35-m automated supernova patrol telescope in Sonoita, AZ. SN 1999bv is located at α = 17^{h}22^{m}40^{s}.73, δ = +60^{o}00′12′′.9 (equinox 2000.0), which is ∼ 2′′.6 west and 24′′.3 south of the nucleus of MCG +10-25-14. The new object was confirmed on a frame taken with the same equipment on Apr. 20.234. The object does not appear on either the first or the second Palomar Sky Survey scans and is absent from archive images taken on 1998 July 13 and 16 using the same equipment. The limiting magnitude on all of the Tenagra I images of this galaxy is ∼ 19.

1999 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green